Consider a fever thermometer that contains 0.01 cm³ of mercury, Find the diameter of the capillary if a 1 C° change corresponds to a level change of 0.5 cm . Assume the glass does noe expand.


1
Expert's answer
2022-02-02T12:12:17-0500

Solution

By changing 1° C change level change l=0.5cm

Volume of mercury is V=0.01 cm³

Volume remains same

Before = after

0.01 cm³=A*0.5

"\frac{\pi D^2}{4}=0.02cm^2"


So diameter D is 0.1596cm.